# Turbulence in stratified rotating topographic wakes

### v1

@@ 4 Parametric dependence of dissipation on stratification and rotation
Besides integrated dissipation, the distribution of instantaneous dissipation and its point-wise correlation to the gradient Richardson is useful for correlating the shear instability to turbulent dissipation. Figure 9 shows the joint probability distribution functions (JPDFs) of the instantaneous total KE dissipation ($\tilde{\varepsilon}$) and the local gradient Richardson number ($Ri_{g}$). Here $\tilde{\varepsilon}$ and $Ri_{g}$ are both based on instantaneous velocity components/buoyancy as previously defined in (9) and (10). The shear-stable background flow with $\tilde{Re}_{b}<0.1$ is excluded from the JPDF calculation. As a consequence, regions of $Ri_{g}>O(1)$ are excluded in the JPDF, allowing a focus on the shear-unstable states. The contour for each case approximates the boundary of the ensemble with $\tilde{Re}_{b}>0.1$ in $(Ri_{g},\tilde{\varepsilon})$ space and is helpful to compare the behavior among cases. Specifically, the contour encloses the JPDF-based top 85% of the ensemble, i.e., the shown contour encloses 85% of the ensemble counting downward in JPDF from its maximum value. By changing the percentage to 80% and 90% similar results are obtained.
The shape of the JPDFs presents noteworthy similarity in shape among all cases. Also, the axis of the JPDF follows
@@
which is a consequence of the scaling analysis in the following paragraph. Equation (14) is plotted in each panel of Figure 9 with the choice of $C_{R}=1/2500$ as the coefficient. Evidently, it is a reasonable approximation to the axis of the JPDF.
Assume a state of stratified layered turbulence where the dissipation is predominantly from the vertical shear of horizontal velocity components, $\tilde{\varepsilon}\sim\nu{S}_{l}^{2}$, where ${S}_{l}^{2}=(\partial_{z}{u})^{2}+(\partial_{z}{v})^{2}$ is the instantaneous, squared shear. The stratification (${N}_{l}$) in the vicinity of the dissipation layers is close to the global constant $N$. The gradient Richardson is hence ${Ri}_{g}=N_{l}^{2}/{S}_{l}^{2}\approx N^{2}/{S}_{l}^{2}$. Thus, the dissipation is ${\tilde{\varepsilon}D}\approx\nu{S}_{l}^{2}\approx\nu N^{2}{Ri_{g}}^{-1}$ and, its inertially-normalized value becomes
$\tilde{\varepsilon}D/U_{\infty}^{3}\approx Fr_{D}^{-2}Re^{-1}{Ri}_{g}^{-1}\,.$ (15)
With $Fr_{D}Re=Re_{N}=900$ and $h/D=0.3$ being constant in the present study, the above relationship further simplifies to (14).
Equation (14) serves as a reference for the instantaneous dissipative regions in the context of stratified turbulence parameters. It can be seen in Fig. 9 that the JPDFs are very narrow around the dashed line given by (14) when $Ri_{g}$ is close to unity and they widen up as $Ri_{g}$ decreases below $1/4$ (more unstable). The low-$Ri_{g}$ spread of $\tilde{\varepsilon}$ is particularly evident at the larger $Fr=0.3$ and $0.4$ values, when the same $Ri_{g}$ corresponds to $\tilde{\varepsilon}$ that ranges in one to two orders of magnitudes and the largest $\tilde{\varepsilon}$ deviates from (14) by an order of magnitude. This is due the intermittency of the local shear and stratification. As strong 3D turbulence is present (large $\tilde{\varepsilon}$), larger local shear and lower local stratification (due to mixing of density) are correlated, both reducing $Ri_{g}$.
@@ 5 Discussion and conclusion
Topographic features are ubiquitous on the seafloor and are hot spots of turbulence generation. Both the physical mechanisms that lead to turbulence and the accurate parametric dependence of dissipation on the overall governing parameters are crucial in the understanding and modeling of bottom ocean flows. To this end, LES of the wake of an isolated 3D topography is employed for a cross-combination of four $Fr$ and five $Ro$, representing moderately strong to strong stratification and rotation rates that range from small MS to small SMS. The LES is conducted at high resolution, sufficient to resolve flow instabilities and the energy-containing scales of wake turbulence. Two instability mechanisms, the KHI and the CI, are found to be the major contributors to turbulent dissipation. Volume-integrated dissipation in the near wake is quantified and rendered in the parameter space $(Fr,Ro)$ and, furthermore, the connection of this parametric dependence to the instability mechanisms is established.
@@
The primary instability in strongly stratified wakes is the KHI between the dislocated layers of velocity–buoyancy structures, albeit oblique rather than horizontal. As stratification increases, the vertical length scale $U_{\infty}/N$ decreases and the distance between the layers becomes thinner (see Fig. 2). The velocity variation over such a distance, largely due to the out-of-phase vortex shedding at different heights, leads to intense shear collocated with the dislocation. The dimensional values of $\varepsilon$ as well as the spiky vertical profiles agree reasonably well with observational measurements. When stratification is relatively weaker, in cases Fr030 and Fr040, the hydraulic response becomes strong, which also breaks down to turbulence via the KHI (see Fig. 2(g,h)).
@@
Rotation influences wake turbulence both indirectly and directly and the latter role depends on stratification. The indirect effect is through the modification of the coherent structures, which was comprehensively studied in Liu et al. (2024). When rotation is weak ($Ro>O(1)$), the vertical structures of wake vortices and dissipation are forward-slanted ‘surfboards’ with $O(U_{\infty}/N)$ vertical thickness, as shown in Fig. 2(a,b), which is a configuration that favors KHI. When rotation is strong ($Ro\leq O(0.1)$), upright ‘columns’ that are reminiscent of stratified Taylor columns are formed (Liu et al., 2024) and the vertical shear and KH turbulence are significantly reduced (see Fig. 2(e)). This indirect effect is also reflected in the volume-integrated dissipation rate for Fr007 and Fr015 cases, shown in Fig. 8, that rotation appears to be solely stabilizing. The direct effect of rotation enters as stratification weakens, in cases Fr030 and Fr040, through the CI. As the vertical length scale increases with increasing $Fr$, its constraint on the CI is released and CI is able to destabilize the flow and change the dependence of dissipation on $Ro$ – a dissipation peak emerges in the SMS range of $Ro=O(0.5)$. The existence of a most destabilizing rotation rate agrees with results from rotating horizontal shear layers (Yanase et al., 1993; Arobone and Sarkar, 2012) and circular flows (Yim et al., 2016), and can be theoretically expected as shown by the linear analysis in Appendix C.
@@
To summarize, the instability mechanisms, KHI and CI, and the effects of stratification and rotation, are strongly intertwined. The KHI results directly from the oblique dislocated layers due to strong stratification, but it is also indirectly influenced by rotation which alters the vertical shear. The CI has a non-monotonic dependence on rotation, while its growth is restricted by stratification. It is the co-existence of KHI and CI, their subsequent nonlinear evolution and cross-dependence/influence that enriches the flow physics, while posting challenges to simple parameterizations. The idealized interaction scheme is shown in Fig. 10.
@@ Acknowledgements.
A similar resolution was found to be able to provide good-quality LES in Liu et al. (2024). The approximate computational cost per case (production run) is listed in Table 2 and the total cost is approximately 2 million CPU hours, not counting the CPU time spent in pre-production scoping simulations.
@@
At sufficiently high Reynolds number, TKE dissipation is expected to be independent of viscosity. In equilibrium turbulence, the universal inertial scaling for dissipation $\langle\varepsilon\rangle=C_{\varepsilon}\mathcal{U}^{3}/\mathcal{L}$ is valid at large Reynolds number limit, with the scaling coefficient being $C_{\varepsilon}\approx 0.1$. Here $\mathcal{U}$ is the fluctuating velocity scale and $\mathcal{L}$ is the integral length scale. At moderate Reynolds number, the ‘constant’ slightly decays as $Re$ increases until the asymptote is reached (Sreenivasan, 1984; Vassilicos, 2015).
[…]
(a) $Fr=0.15,Ro=0.15$
(b) $Fr=0.15,Ro=0.75$
(c) $Fr=0.15,Ro=1.5$
@@
[…]
@@
The starting point is the non-dimensional, linearized vorticity equation
@@
where primes denote perturbations with respect to the base flow (denoted by brackets). The operator $\bar{D}/\bar{D}t=\partial_{t}+\langle u_{i}\rangle\partial_{x_{i}}$ represents mean convection and $S_{ij}$ is the rate-of-strain tensor. The free indices $i=1,2,3$ denote the three spatial dimensions that are $(x,y,z)$ in Cartesian coordinates and are $(r,\theta,z)$ in cylindrical coordinates, and the corresponding velocity components are $(u,v,w)$. The base flow is $\langle u\rangle=U_{SL}(y)$ in the parallel shear flow case and is $\langle v\rangle=\langle v\rangle(r)$ in the axisymmetric case. The buoyancy Richardson number is $Ri_{b}=Fr_{h}^{-2}$, where $Fr_{h}$ is the horizontal Froude number, and the non-dimensional Coriolis parameter is $f_{c}=Ro^{-1}$. Viscous effects are neglected ($Ek$?).
@@
$-f_{c}(\langle\omega_{z}\rangle+f_{c})>0,$ (21)
there is a pair of real eigenvalues of opposite signs, with the positive one $\lambda_{+}=[-f_{c}(f_{c}+\omega_{z})]^{1/2}$ corresponding to instability. Otherwise, a pair of purely imaginary, conjugate eigenvalues will imply inertial waves.
The condition (21) is equivalent to the absolute vorticity criterion for inertial instability (Holton, 1972). In particular, for a base flow with shear $\langle\omega_{z}\rangle$, the growth rate of the perturbations is the largest when $f_{c}=f_{c,{\rm max}}=-\langle\omega_{z}\rangle/2$, equivalently $Ro_{{\rm max}}=\langle\omega_{z}\rangle/f_{c,{\rm max}}=-2$ (anticyclonic). In other words, the most destabilizing (anticyclonic) rotation with respect to the given shear, is in the SMS range. When rotation is very strong ($Ro\ll O(1)$), $|f_{c}|\gg|\langle\omega_{z}\rangle|$ and hence $-f_{c}(\langle\omega_{z}\rangle+f_{c})\approx-f_{c}^{2}<0$, the CI mode is stabilized, consistent with Yanase et al. (1993) and Arobone and Sarkar (2012). When rotation is very weak ($Ro\gg O(1)$), the former analysis does not apply. We also note that the optimal growth rate may not always be achieved, but nevertheless, it serves as a good predictor for the strength of the CI.
The stability of the system (19) can be interpreted as the response of the shear layer to coordinate rotation. The system is the most resonant when the forcing frequency ($f_{c}$) is the closest to the intrinsic frequency of the system ($\langle\omega_{z}\rangle/2$, the angular velocity of the solid-body rotation part of fluid motions). The amplification mechanism is through the linearized Navier–Stokes.
@@
$-\chi=-\left(\frac{2\langle v\rangle}{r}+f_{c}\right)(\langle\omega_{z}\rangle+f_{c})>0,$ (28)
where $\chi$ is exactly the generalized Rayleigh discriminant (Kloosterziel and Van Heijst, 1991; Mutabazi et al., 1992). Taking $\lambda_{+}(r)=\sqrt{-\chi(r)}$ as the estimated local growth rate for regions with $\chi<0$, the most destabilizing rotation rate can be searched for any axisymmetric eddy profile. It is typical that an annular region satisfies $\chi<0$ and becomes unstable. Taking the derivative of (27) with respect to $f_{c}$, the locally most destabilizing rotation rate is $f_{c,{\rm max}}(r)=-(\langle v\rangle/r+\langle\omega_{z}\rangle/2)$. Although it should be evaluated with a global measure and it is flow-specific, the formal expression of $f_{c,{\rm max}}$ still suggests that an intermediate rotation rate at $Ro=O(1)$ will lead to the fastest three-dimensionalization.

### v2

@@ 4.2 Relating dissipation rate to local stability criteria
Figure 10 shows the joint probability distribution functions (JPDFs) of the instantaneous total KE dissipation ($\tilde{\varepsilon}$) and the local gradient Richardson number ($Ri_{g}$). Here $\tilde{\varepsilon}$ and $Ri_{g}$ are both based on instantaneous velocity components/buoyancy as previously defined in (9) and (10). The shear-stable background flow with $\tilde{Re}_{b}<0.1,Ri_{g}>O(1)$ is excluded from the JPDF, allowing a focus on the shear-unstable states. The contours in each plot enclose the most probable 85% of the data/ensemble. When this percentage is changed to 80% and 90%, similar results are obtained. The shape of the JPDFs presents noteworthy similarity in shape among all cases. Also, the axis of the JPDF follows
@@
plotted as dashed lines in Fig. 10. Evidently, (14) provides a quantitative relationship – a power law – between dissipation and $Ri_{g}$ across a range of stable and unstable $Ri_{g}$ and at different $Fr$. The following scaling analysis is useful to deduce (14).
In the limit of stable, strongly stratified flow, dissipation is predominantly from the vertical shear of horizontal velocity components, $\tilde{\varepsilon}\approx\nu{S}_{l}^{2}$, where ${S}_{l}^{2}=(\partial_{z}{u})^{2}+(\partial_{z}{v})^{2}$ is the instantaneous, squared shear. The local stratification (${N}_{l}$) is close to the global constant $N$, leading to a gradient Richardson number of ${Ri}_{g}=N_{l}^{2}/{S}_{l}^{2}\approx N^{2}/{S}_{l}^{2}$. The dissipation can then be expressed as ${\tilde{\varepsilon}}\approx\nu{S}_{l}^{2}\approx\nu N^{2}{Ri_{g}}^{-1}$. In another limit of fully turbulent, nearly isotropic turbulence, $\tilde{\varepsilon}=\nu(\partial_{j}u_{i})^{2}\approx\frac{15}{2}\nu(\partial_{z}u)^{2}=\frac{15}{4}\nu S_{l}^{2}$, invoking isotropic statistical relations. Although the proportionality constant is different in the two limits, the difference is small on a logarithmic scale and the power-law dependence remains the same. The inertially-normalized relation becomes
$\tilde{\varepsilon}D/U_{\infty}^{3}\approx Fr_{D}^{-2}Re^{-1}{Ri}_{g}^{-1},$ (15)
which simplifies to (14) with $Fr_{D}Re=Re_{N}=900$ being a constant and $Fr_{D}=Fr(h/D)$ being linearly proportional to the case-dependent $Fr$.
Equation (14) serves as a reference for the expected instantaneous dissipation in the context of stratified turbulence parameters. It can be seen in Fig. 10 that the JPDFs are very narrow around the dashed line given by (14) when $Ri_{g}$ is close to unity, corresponding to the stable limit. As $Ri_{g}$ decreases and the transition to instability occurs, the JPDFs widen as there is scatter in dissipation. The low-$Ri_{g}$ spread of the JPDF of $\tilde{\varepsilon}$ is particularly evident at the larger $Fr=0.3$ and $0.4$ values, where the same $Ri_{g}$ corresponds to $\tilde{\varepsilon}$ that differ by one to two orders of magnitude. This is due to the intermittency of the local shear and stratification. As strong 3D turbulence is present, both larger local shear and lower local stratification (due to mixing of density) occur, and it becomes causally unclear whether lower $Ri_{g}$ leads to turbulence or turbulent overturning motions reduces $Ri_{g}$. But in all, (14) serves as a reasonable, quantitative prediction of the turbulent dissipation provided the local index $Ri_{g}$.
On the other hand, it is more challenging to predict dissipation using the local PV. The JPDFs of local dissipation and PV (not shown) display poor correlation – given one, there is a wide scatter in the distribution of another. This is simply due to the fact that fully-developed turbulence contains intense vorticity and hence large PV (of both signs). While large-scale PV in the base flow may imply (in)stability, small-scale PV in fully turbulent flow is the result rather than the cause of turbulence.
@@ 5 Discussion and conclusion
Topographic features are ubiquitous on the seafloor and are hot spots of turbulence generation. Both the physical mechanisms that lead to turbulence and the accurate parametric dependence of dissipation on the overall governing parameters are crucial in the understanding and modeling of bottom ocean flows. To this end, LES of the wake of an isolated 3D topography is employed for a cross-combination of four $Fr$ and five $Ro$, representing moderately strong to strong stratification and rotation rates that range from small MS to small SMS. The LES is conducted at high resolution, sufficient to resolve flow instabilities and the energy-containing scales of wake turbulence. Our results indicate that turbulence and dissipation in the near wake stem from two instabilities, KHI and CI, and the case-dependent intensity of dissipation depends on the development and manifestation of these instabilities. Volume-integrated dissipation in the near wake is quantified and rendered in the $(Fr,Ro)$ parameter space and, furthermore, the connection of this parametric dependence to the instability mechanisms is established.
@@
The primary instability in strongly stratified wakes is KHI between the dislocated layers of velocity–buoyancy structures, which are oblique rather than horizontal. As stratification increases, the vertical length scale $U_{\infty}/N$ decreases and the distance between the layers becomes smaller (see Fig. 3). The velocity variation over such a distance, largely due to the out-of-phase vortex shedding at different heights, leads to intense shear collocated with the dislocation. The dimensional values of $\varepsilon$ as well as the spiky vertical profiles agree reasonably well with observational measurements. When stratification is relatively weaker, in cases Fr030 and Fr040, the hydraulic response becomes strong, which also breaks down to turbulence via KHI (see Fig. 3(g,h)).
@@
Rotation influences wake turbulence both indirectly and directly and the latter role depends on stratification. The indirect effect is through the modification of the coherent structures, which was comprehensively studied in Liu et al. (2024). When rotation is weak ($Ro>O(1)$), the vertical structures of wake vortices and dissipation are forward-slanted ‘surfboards’ with $O(U_{\infty}/N)$ vertical thickness, as shown in Fig. 3(a,b), which is a configuration that favors KHI. When rotation is strong ($Ro\leq O(0.1)$), upright ‘columns’ that are reminiscent of stratified Taylor columns are formed (Liu et al., 2024) and the vertical shear and KH turbulence are significantly reduced (see Fig. 3(e)). This indirect effect is also reflected in the volume-integrated dissipation rate for Fr007 and Fr015 cases, shown in Fig. 9, that rotation appears to be solely stabilizing. The direct effect of rotation enters as stratification weakens, in cases Fr030 and Fr040, through CI. As the vertical length scale increases with increasing $Fr$, its constraint on CI is released and CI is able to destabilize the flow and change the dependence of dissipation on $Ro$ – a dissipation peak emerges in the SMS range of $Ro=O(0.5)$. The existence of a most destabilizing rotation rate agrees with results from rotating horizontal shear layers (Yanase et al., 1993; Arobone and Sarkar, 2012) and circular flows (Yim et al., 2016), and can be theoretically expected as shown by the linear analysis in Appendix C.
@@
To summarize, the instability mechanisms, KHI and CI, and the effects of stratification and rotation, are strongly intertwined. The KHI results directly from the oblique dislocated layers due to strong stratification, but it is also indirectly influenced by rotation which alters the vertical shear. The CI has a non-monotonic dependence on rotation, while its growth is restricted by stratification. It is the co-existence of KHI and CI, their subsequent nonlinear evolution and cross-dependence/influence that enriches the flow physics, while posing challenges to simple parameterizations. The idealized interaction between stratification and rotation in the context of wake turbulence is shown as a schematic in Fig. 11.
@@ Acknowledgements.
A similar resolution was found to provide good-quality LES in Liu et al. (2024). The approximate computational cost per case (production run) is listed in Table 2 and the total cost is approximately 2 million CPU hours, not counting the CPU time spent in pre-production scoping simulations.
@@
At sufficiently high Reynolds number, TKE dissipation is expected to be independent of viscosity. In equilibrium turbulence, the universal inertial scaling for dissipation $\langle\varepsilon\rangle=C_{\varepsilon}\mathcal{U}^{3}/\mathcal{L}$ is valid at large Reynolds number limit. Here $\mathcal{U}$ is the fluctuating velocity scale and $\mathcal{L}$ is the integral length scale. At moderate Reynolds number, the ‘constant’ slightly decays as $Re$ increases until the asymptote is reached (Sreenivasan, 1984; Vassilicos, 2015).
In order to examine the sensitivity of our results to the Reynolds number, three cases, Fr015Ro015, Fr015Ro075, Fr015Ro1p5, are selected for the $Re$-sensitivity study. They are run at Re=20  [^0] Re=20\,000 according to Table 1. Simulations at two additional Reynolds numbers, Re=10 000,Re=30  [^0] Re=10\,000,\,Re=30\,000, are conducted for each of the three cases.
@@
Figure 13: Similar to Fig. 3. The Reynolds numbers are Re=10  [^0] Re=10\,000 ($Re_{N}=450$) in (a,b) and Re=30  [^0] Re=30\,000 ($Re_{N}=1350$) in (c,d). Panels (c,d) look very similar to Fig. 3 (c-d), which are at Re=20  [^0] Re=20\,000 ($Re_{N}=900$).
@@
The starting point is the non-dimensional, linearized vorticity equation for inviscid perturbations:
@@
obtained by taking the curl of the perturbed momentum equations linearized around a base flow. The base flow is denoted by brackets and the perturbations by primes. The operator $\bar{D}/\bar{D}t=\partial_{t}+\langle u_{i}\rangle\partial_{x_{i}}$ represents mean convection and $S_{ij}$ is the rate-of-strain tensor. The free indices $i=1,2,3$ denote the three spatial dimensions that are $(x,y,z)$ in Cartesian coordinates and $(r,\theta,z)$ in cylindrical coordinates. The corresponding velocity components are $(u,v,w)$. The base flow is $\langle u\rangle=U_{SL}(y)$ in the parallel flow case and is $\langle v\rangle=\langle v\rangle(r)=\langle u_{\theta}\rangle(r)$ in the axisymmetric flow case. The buoyancy Richardson number is $Ri_{b}=Fr_{h}^{-2}$, where $Fr_{h}$ is the horizontal Froude number, and the non-dimensional Coriolis parameter is $f_{c}=Ro^{-1}$.
@@
$f_{c}(\langle\omega_{z}\rangle+f_{c})<0,$ (21)
there is a pair of real eigenvalues of opposite signs, with the positive one $\lambda_{+}=\sqrt{-f_{c}(f_{c}+\langle\omega_{z}\rangle)}$ corresponding to instability. Otherwise, a pair of purely imaginary, conjugate eigenvalues will imply inertial waves.
The condition (21) is equivalent to the absolute vorticity criterion for inertial instability (Holton, 1972). In particular, for a base flow with shear $\langle\omega_{z}\rangle$, the growth rate of the perturbations is the largest when $f_{c}=f_{c,{\rm max}}=-\langle\omega_{z}\rangle/2$, equivalently $Ro_{{\rm max}}=\langle\omega_{z}\rangle/f_{c,{\rm max}}=-2$ (anticyclonic). In other words, the most destabilizing rotation is anticyclonic (with respect to the given shear), and is in the SMS range. Cyclonic shear ($f_{c}$ with the same sign as $\langle\omega_{z}\rangle$) always result in stability according to (20). When rotation is very strong compared to the shear ($Ro\ll O(1)$, $|f_{c}|\gg|\langle\omega_{z}\rangle|$), $\lambda^{2}=-f_{c}(\langle\omega_{z}\rangle+f_{c})\approx-f_{c}^{2}<0$, the CI mode is stabilized regardless of the sign of $f_{c}$, consistent with Yanase et al. (1993) and Arobone and Sarkar (2012). When rotation is very weak ($Ro\gg O(1)$, $f_{c}\approx 0$), the predicted growth rate (20) vanishes since $f_{c}$ is small and $(\langle\omega_{z}\rangle+f_{c})$ is finite. Thus, CI is stabilized at either strong or weak rotation. We also note that the optimal growth rate may not always be achieved, but nevertheless, it serves as a good predictor for the strength of CI.
The stability of the system (19) can be interpreted as the response of the shear layer to coordinate rotation. The system is the most resonant when the forcing frequency ($f_{c}$) is the closest to the intrinsic frequency of the system ($\langle\omega_{z}\rangle/2$, the angular velocity of the solid-body rotation part of fluid motions). The initial amplification mechanism is linear.
@@
$\chi=\left(\frac{2\langle v\rangle}{r}+f_{c}\right)(\langle\omega_{z}\rangle+f_{c})<0,$ (28)
where $\chi$ is exactly the generalized Rayleigh discriminant (Kloosterziel and Van Heijst, 1991; Mutabazi et al., 1992). Taking $\lambda_{+}(r)=\sqrt{-\chi(r)}$ as the estimated local growth rate for regions with $\chi<0$, the most destabilizing rotation rate can be searched for any axisymmetric eddy profile. It is typical that an annular region satisfies $\chi<0$ and becomes unstable. Taking the derivative of (27) with respect to $f_{c}$, the locally most destabilizing rotation rate is $f_{c,{\rm max}}(r)=-(\langle v\rangle/r+\langle\omega_{z}\rangle/2)$. Although it should be evaluated with a global measure and it is flow-specific, the formal expression of $f_{c,{\rm max}}$ still suggests that an intermediate rotation rate at $Ro=O(1)$ will lead to the fastest three-dimensionalization. It is also noted that, unlike in parallel flows, for which the instability criterion (21) can only be met with anticyclonic shear ($\langle\omega_{z}\rangle$ and $f_{c}$ have opposite signs), in axisymmetric flows both cyclonic and anticyclonic vortices can become unstable to CI.
The implication of this Appendix is that the criteria (21) and (28) are not only useful as stability discriminators, but they are also connected to how fast the perturbations can grow. The latter fact is utilized to predict the most destabilizing rotation rate for a given base flow, and has led to the understanding that intermediate rotation rates are the most unstable for CI, theoretically explaining the findings in section 4.